satellite73 (satellite73):

It takes Harland 15 minutes to rake all leaves in the front yard. If Trudy scatters the leaves while Harland rakes, it takes him 20 minutes to rake the leaves. How many minutes does it take Trudy to scatter all the leaves?

OpenStudy (anonymous):

just posted because this has to be the dumbest question i have ever seen

ganeshie8 (ganeshie8):

lol they're quantifying the scattering

OpenStudy (anonymous):

i tried to help, came up with the math teacher answer of 60, but god knows if i am right i figured get a leaf blower and scatter them is two seconds

ganeshie8 (ganeshie8):

\[\dfrac{1}{15}+\dfrac{1}{x} = \dfrac{1}{20}\]

OpenStudy (anonymous):

man now there is another one!! this teacher needs some help!

ganeshie8 (ganeshie8):

the teacher is hilarious lmao

OpenStudy (anonymous):

part 2 Perry can fill a sorter with oranges from the conveyer belt in 10 minutes. While Perry fills the sorter, Gillian takes oranges out of the sorter and puts them in shipping crates. With Gillian taking oranges out of the sorter, it takes 25 minutes for Perry to fill the sorter.

OpenStudy (anonymous):

lolololol!